# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### 14 Oct 2024 v1

@@ Appendix B File Sharing Platforms
Appendix C Deepfake production guides we examined
@@
Appendix D Breakdown of sexual deepfake video-related discussion topics
Total: 6,259 threads comprising 33,496 posts in the Adult DeepFake Videos (NSFW / 18+) and DeepFake Video Creation Tools sections.
@@ Appendix E Paid Request Codebook
Code Meaning
pleasure sexual satisfaction, fantasy, or preference
bullying repeated psychological harm done to victim
circum_consent circumvent the need to obtain consent
revenge counteraction to cause harm to victim in return for an "injury"
masculinity proving masculinity to a sexually deviant network (Henry & Flynn 2019)
@@
sextortion forcing a victim to do (sexual) acts by threatening share their explicit images
reputation cause injury to social reputation of victim
community collective archive for a community to enjoy or benefit from
private_use requester specifies they want this video for private use
@@
join_reason requester implies the request is what brought them to the forum
fetish fetish content as defined in DSM-5
art discussion of content as artistic or the deepfake creators as artists
@@
delusional discussion that positions the requester/payer as entitled to sexualizing the target
@@
show_skill request framing as an opportunity for creator to display skill/prowess over others
@@
creation_interest requester expresses interest in learning how to create their own deepfakes
sfw requester wants a SFW video
lf_csam requester is seeking CSAM
lf_model requester is seeking a model and not the output video
lf_faceset requester is seeking a faceset (image dataset)
@@
lf_data_labor requester is seeking labor for making deepfake image data usable
lf_audio requester is seeking audio fake (replacement of voice)
@@
batch_request requesting multiple videos (can imply a long chain of commissions)
@@
provides_src requester describes private source materials that they can provide
closed thread closed for unknown reason
completed thread closed because the request was fulfilled
has_price includes some price figure (concrete) or rate
pm offers a private communication channel
quality discussion of desire of good/high-quality material
@@
professional discussion of desire of professionalism in their chosen deepfake creator
turnaround discussion of a concrete turnaround time from creator
ethics requester discusses the ethical considerations in some capacity
legality requester discusses the legality in some capacity
compliance requester discusses the compliance with the community rules
extra_multi_target willing to pay extra for multiple targets
extra_male willing to pay extra because target is male
extra_angle willing to pay extra because the angle is challenging
@@
extra_body willing to pay extra or wait longer because of challenges posed by target body match
target_sw specified target already produces sexual content
target_obfus if the target’s name is not specified anywhere in the request
target_personal specified target is from the requester’s personal life
target_self specified target is the requester themself
trust_verified requester is seeking a verified creator (designation of reputation on MDF)
@@
trust_reviews requester is seeking a creator that has a lot of reviews from prior commissions
@@
trust_portfolio requester is seeking portfolio of candidate creators’ past work to vet them
@@
Table 5: Codebook for need-finding interviews —Codebook for categorizing the different themes that emerged from our need-finding interviews with eight journalists.

### 02 Jan 2025 v2

@@ Appendix B File Sharing Platforms
Appendix C Investigated Deepfake production guides we examinedProduction Guides
The full list of the 25 guides and tutorials forum posts we examined for manual content analysis.
@@ Appendix C Investigated Deepfake production guides we examinedProduction Guides
Appendix D Breakdown of sexual Sexual deepfake video-related video discussiontopics
Total: The breakdown, by number of threads and posts, of forum discussion topics relating to sexual deepfake videos. The total number of threads is 6,259threads , comprising 33,496 posts in the Adult DeepFake Videos (NSFW / 18+) and DeepFake Video Creation Tools sections.
@@ Appendix E Paid Request Codebook
Code Meaning %
pleasure sexual satisfaction, fantasy, or preference 23.7
bullying repeated psychological harm done to victim 0
circum_consent circumvent the need to obtain consent 0
revenge counteraction to cause harm to victim in return for an "injury" 0
masculinity proving masculinity to a sexually deviant network (Henry & Flynn 2019) 0.8
@@
sextortion forcing a victim to do (sexual) acts by threatening share their explicit images 0
reputation cause injury to social reputation of victim 0
community collective archive for a community to enjoy or benefit from 7.1
private_use requester specifies they want this video for private use 1.9
@@
join_reason requester implies the request is what brought them to the forum 2.6
fetish fetish content as defined in DSM-5 1.1
art discussion of content as artistic or the deepfake creators as artists 2.6
@@
delusional discussion that positions the requester/payer as entitled to sexualizing the target 0.8
@@
show_skill request framing as an opportunity for creator to display skill/prowess over others 0.8
@@
creation_interest requester expresses interest in learning how to create their own deepfakes 1.5
sfw requester wants a SFW video 1.5
lf_csam requester is seeking CSAM 0.4
lf_model requester is seeking a model and not the output video 1.1
lf_faceset requester is seeking a faceset (image dataset) 0.4
@@
lf_data_labor requester is seeking labor for making deepfake image data usable 0.4
lf_audio requester is seeking audio fake (replacement of voice) 1.1
@@
batch_request requesting multiple videos (can imply a long chain of commissions) 4.9
@@
provides_src requester describes private source materials that they can provide 11.7
closed thread closed for unknown reason 1.5
completed thread closed because the request was fulfilled 0.8
has_price includes some price figure (concrete) or rate 4.1
pm offers a private communication channel 15.0
quality discussion of desire of good/high-quality material 12.0
@@
professional discussion of desire of professionalism in their chosen deepfake creator 0.4
turnaround discussion of a concrete turnaround time from creator 1.9
ethics requester discusses the ethical considerations in some capacity 1.5
legality requester discusses the legality in some capacity 0.8
compliance requester discusses the compliance with the community rules 0.8
extra_multi_target willing to pay extra for multiple targets 0.4
extra_male willing to pay extra because target is male 0.4
extra_angle willing to pay extra because the angle is challenging 0.4
@@
extra_body willing to pay extra or wait longer because of challenges posed by target body match 0.8
target_sw specified target already produces sexual content 1.5
target_obfus if the target’s name is not specified anywhere in the request 10.9
target_personal specified target is from the requester’s personal life 2.3
target_self specified target is the requester themself 0.8
trust_verified requester is seeking a verified creator (designation of reputation on MDF) 0.4
@@
trust_reviews requester is seeking a creator that has a lot of reviews from prior commissions 0.4
@@
trust_portfolio requester is seeking portfolio of candidate creators’ past work to vet them 2.6
@@
Table 5: Codebook for need-finding interviews Codebook for Paid Deepfake Requests —Codebook for categorizing the different themes that emerged from our need-finding interviews with eight journaliststhematic analysis of a sample (N=263) of paid requests for deepfake videos.